OpenStudy (anonymous):

Sorry for asking so many questions but I'm completely lost so here's another one. How would I simplify this problem and rewrite it in an equivalent form with positive exponents (-x^-3)^-5

OpenStudy (anonymous):

\[a^{-2}=1/a^2\]

OpenStudy (jlastino):

Just make it 1/(given equation) (because it is raised to a negative exponent)

OpenStudy (anonymous):

So I would make it a division problem insted of a multiplication?

OpenStudy (jlastino):

I think so but what you are doing is just expressing it more clearly (simplifying)

OpenStudy (anonymous):

Would I add the two exponents together or just leave it the same? 1/(-x^-3)^-5, like that

OpenStudy (jlastino):

You make the exponent positive

OpenStudy (jlastino):

then multiply since (x^a)^b = (x^ab)

OpenStudy (anonymous):

Ok, I think I understand now. Thanks for the help :)

OpenStudy (jlastino):

Could you show me the answer? :D

OpenStudy (anonymous):

(-x^-3)^-5= (-x^8), is that corret?

OpenStudy (jlastino):

No....keep in mind when you raise an exponent by an exponent you multiply them

OpenStudy (anonymous):

Oh so I would multiply -3 by -5

OpenStudy (anonymous):

Is the correct answer 1/(-x^15)

OpenStudy (jlastino):

i think so, yeah

OpenStudy (jlastino):

oh wait...it should be -x^15

OpenStudy (anonymous):

Ok, thanks